OpenStudy (anonymous):

Find the difference. 4ab2 - 3ab2

OpenStudy (anonymous):

1ab2 they are like terms because they have the same variables "ab2" so you just subtract the coeficients

OpenStudy (skullpatrol):

\[4ab ^{2 } -3ab ^{2} = (4-3)ab ^{2}\]

OpenStudy (anonymous):

well what would the answer be? :O

OpenStudy (anonymous):

1ab2

OpenStudy (anonymous):

just subtract 4-3

OpenStudy (anonymous):

so its just 1?

OpenStudy (anonymous):

yah

OpenStudy (skullpatrol):

Have you ever seen a(b + c) = ab + ac

OpenStudy (anonymous):

well no its 1ab2

OpenStudy (anonymous):

if the variables are the same( "ab2") then u just subtrac the first numbers then you add the variables back at the end

OpenStudy (anonymous):

never seen any of this its a pretest

OpenStudy (anonymous):

never seen any of this its a pretest

OpenStudy (skullpatrol):